OpenStudy (anonymous):

One month, a music site observed that 60% of the people who downloaded songs from its site downloaded Q Sam's latest single. The equation below represents this information, where x represents the total number of people who downloaded songs from the site that month. x = 0.6x + 384 How many people who downloaded songs from the site that month downloaded Q Sam's latest single? 576 240 144 960

OpenStudy (smored):

all you have to do is solve for x. \[x=0.6x+384\]subtract 0.6x from both sides\[0.4x=384\]divide both sides by 0.4\[x=960\]

OpenStudy (anonymous):

Thank you hun :)

OpenStudy (smored):

\[\frac{ x }{ 0.4 }=x \div(\frac{ 4 }{ 10 })=x(\frac{ 10 }{ 4 })=x(2.5)\]